OpenStudy (anonymous):

P(eight, non face card) in 52 deck

OpenStudy (anonymous):

whats the comma mean?

OpenStudy (anonymous):

probability that it is an eight or a non face card

OpenStudy (anonymous):

so do you know what operation we do when there is an OR?

OpenStudy (anonymous):

p(a or b) = p(a) + p(b) - p(a u b)

OpenStudy (anonymous):

okay so it would be (4/52+ 44/52) -4/52 = 11/13?

OpenStudy (anonymous):

seems right